# Bandit and Delayed Feedback in Online Structured Prediction

### v1

@@ D.1 Concentration inequality
where we recall $C_{x}=\mathrm{diam}(\mathcal{X})$. From this inequality,
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$ $\leq 2\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2C_{x}^{2}\mathbb{E}_{t}\left[{\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\leq 2bS_{t}(\bm{W}_{t})+2C_{x}^{2}\left(\|\bm{y}_{t}\|_{2}^{2}-2\bm{y}_{t}^{\top}\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]+\mathbb{E}_{t}\left[{\|\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\right)$
@@
Finally, we are ready to prove Theorem 3.8.
Proof of Theorem 3.8.
@@
where the first inequality follows from Lemma 3.3 and the last inequality follows from Lemma D.6 and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, by combining with the last inequality, we have
@@
Here, we derive the regret upper bounds provided by the algorithm established in Theorem 3.8 for online multiclass classification, online multilabel classification, and ranking. Recall that we can achieve
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(\omega^{1/3}\left(BC_{x}T\right)^{2/3}\right),$ (11)
where we recall that $\omega$ is defined as $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$. Note that when $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, then the matrix $\bm{Q}$ is invertible and $\lambda_{\min}(\bm{Q})>0$, and thus
$\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)=\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\leq\frac{1}{\lambda_{\min}(\bm{Q})}\sum_{i=1}^{d}\|(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\|_{2}^{2}\leq\frac{1}{\lambda_{\min}(\bm{Q})}\|\bm{V}^{-1}\|_{\mathrm{F}}^{2}.$ (12)
In each problem setting, this regret upper bound can be reduced to the following bounds:
@@
We first consider multiclass classification with the 0-1 loss. Since $\bm{V}=\bm{1}\bm{1}^{\top}-\bm{I}$, we have $\|{\bm{V}}^{-1}\|_{\mathrm{F}}^{2}\leq d$ for $d\geq 2$. Recalling that $\mu$ is the uniform distribution over $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$, we have $\mathbb{E}_{\bm{y}\sim\mu}[(\bm{y}^{\top}\bm{x})^{2}]=\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}$ for any $\bm{x}\in\mathbb{R}^{d}$. Hence, $\lambda_{\min}(\bm{Q})=\min_{\|\bm{x}\|_{2}=1}\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]=\frac{1}{d}$, where the first equality is from Cesa-Bianchi & Lugosi (2012, Lemma 2). Since $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$ in this case, from (12) we can let $\omega=d/\lambda_{\min}(\bm{Q})=d^{2}$. Substituting these into our upper bound in (11), we obtain
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((BC_{x}dT)^{2/3}\right).$
@@
Hence, we have $\lambda_{\min}(\bm{Q})=\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]\geq\frac{m(d-m)}{d^{2}}$, where the equality is from Cesa-Bianchi & Lugosi (2012, Lemma 2). Since $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, from (12) we can choose $\omega=d^{3}/(4\lambda_{\min}(\bm{Q}))=\frac{d^{5}}{4m(d-m)}$. Therefore, our regret upper bound in (11) is reduced to
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(\left(\frac{B^{2}C_{x}^{2}d^{5}}{m(d-m)}\right)^{1/3}T^{2/3}\right).$
@@
We finally consider online ranking with the Hamming loss and the number of items $m$. From Cesa-Bianchi & Lugosi (2012, Proposition 4), the smallest positive eigenvalue is at least $1/m$. Hence, since $\bm{V}=-\frac{1}{m}\bm{I}$, we have
@@
where we used $\mathrm{rank}(\bm{Q}^{+})\leq d=m^{2}$, and this allows us to choose $\omega=m^{5}$. Substituting these values into our regret upper bound in (11) , we obtain
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(m^{5/3}(BC_{x}T)^{2/3}\right).$
Appendix E Omitted Details of Section 4
@@ E.1 Details of Optimistic Delayed Adaptive FTRL (ODAFTRL)
We provide a more detailed explanation of the Optimistic Delayed Adaptive FTRL (ODAFTRL) algorithm used for updating $\bm{W}_{t}$ in Section 4. We recall that ODAFTRL computes $\bm{W}_{t}$ by the following update rule:
@@
where $\lambda_{t}\geq 0$ is a regularization parameter. The ODAFTRL algorithm, when using this parameter update called AdaHedgeD, satisfies the following lemma:
Lemma E.1 (Flaspohler et al. 2021, Theorem 12).
Fix $\alpha>0$. Let $S_{t}:\mathcal{W}\to\mathbb{R}$ be a convex function for each $t=1,\dots,T$. Suppose that we update $\lambda_{t}$ in (13) by the following AdaHedgeD update:
@@
$\delta_{t}$ $=\min\left\{F_{t+1}(\bm{W}_{t})-F_{t+1}(\bar{\bm{W}}_{t}),\langle\bm{G}_{t},\bm{W}_{t}-\bar{\bm{W}}_{t}\rangle,F_{t+1}(\widehat{\bm{W}}_{t})-F_{t+1}(\bar{\bm{W}}_{t})+\langle\bm{G}_{t},\bm{W}_{t}-\widehat{\bm{W}}_{t}\rangle\right\}_{+},$
@@
$\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\leq\sum_{t=1}^{T}{S_{t}(\bm{U})}+\left(\frac{B^{2}}{2\alpha}+1\right)\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+2\alpha b_{t}}\right),$
@@
In the following, we let $\alpha=\max_{\bm{W}\in\mathcal{W}}\frac{\|\bm{W}\|_{\mathrm{F}}^{2}}{2}=\frac{B^{2}}{2}$ for simplicity. Then,
$\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\leq\sum_{t=1}^{T}{S_{t}(\bm{U})}+2\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+B^{2}b_{t}}\right),$
@@
We present Theorem 4.3 in a more detailed form and provide its proof.
@@
Let $\alpha=\frac{B^{2}}{2}$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a delay of $D$ achieves
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BDL+\frac{2bB^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3}.$
@@
From Lemma E.1 and the definition of $a_{t}$ and $b_{t}$, we have
$\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}$ $\leq 2\left(2B\max_{s\in[T]}\sum_{i=s-D}^{s-1}\|\bm{G}_{i}\|_{\mathrm{F}}+\sqrt{\sum_{t=1}^{T}{\left(B^{2}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+B^{2}\|\bm{G}_{t-D:t}\|_{\mathrm{F}}\|\bm{G}_{t}\|_{\mathrm{F}}\right)}}\right)$
$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right),$
where we used the Cauchy–Schwarz inequality, $\|\bm{G}_{t}\|_{\mathrm{F}}\leq L$, $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$, and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$ in the last inequality. From this inequality and Assumption 4.1, we can evaluate surrogate regret as
@@
$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right)-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
$\leq 4BDL+\frac{2bB^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3}.$
We used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq({3}/{4})\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$ that hold for any $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$ in the last inequality. ∎
E.3 Proof of Theorem 4.4
We present Theorem 4.4 in a more detailed form and provide its proof.

### v2

@@ D.2 Proof of high probability bound
where we recall $C_{x}=\mathrm{diam}(\mathcal{X})$. From this inequality, we obtain
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$ $\leq 2\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2C_{x}^{2}\mathbb{E}_{t}\left[{\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]$
$\leq 2bS_{t}(\bm{W}_{t})+2C_{x}^{2}\left(\|\bm{y}_{t}\|_{2}^{2}-2\bm{y}_{t}^{\top}\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]+\mathbb{E}_{t}\left[{\|\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\right)$
@@
Finally, we are ready to prove Theorem D.4.
Proof of Theorem D.4.
@@
where the first inequality follows from Lemma 3.3 and the last inequality follows from Lemma D.7 and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, we obtain
@@
We derive the surrogate regret upper bounds provided by the algorithm established in Theorem D.4 for online multiclass classification, online multilabel classification, and ranking. Recall that we can achieve
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2^{5/3}\omega^{1/3}\left(BC_{x}T\right)^{2/3},$ (8)
where we recall that $\omega$ is defined as $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$. Note that when $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, then the matrix $\bm{Q}$ is invertible and $\lambda_{\min}(\bm{Q})>0$, and hence
$\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)$ $=\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\leq\frac{1}{\lambda_{\min}(\bm{Q})}\sum_{i=1}^{d}\|(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\|_{2}^{2}$
$\leq\frac{1}{\lambda_{\min}(\bm{Q})}\|\bm{V}^{-1}\|_{\mathrm{F}}^{2}.$ (9)
Consequently, surrogate regret bounds for specific problems are obtained as follows:
@@
We first consider multiclass classification with the 0-1 loss. From $\bm{V}=\bm{1}\bm{1}^{\top}-\bm{I}$, we have $\|{\bm{V}}^{-1}\|_{\mathrm{F}}^{2}\leq d$ for $d\geq 2$. Recalling that $\mu$ is the uniform distribution over $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$, we have $\mathbb{E}_{\bm{y}\sim\mu}[(\bm{y}^{\top}\bm{x})^{2}]=\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}$ for any $\bm{x}\in\mathbb{R}^{d}$. Hence, we have $\lambda_{\min}(\bm{Q})=\min_{\|\bm{x}\|_{2}=1}\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]=\frac{1}{d}$, where the first equality follows from [8, Lemma 2]. Since $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$ holds in this case, from (D.4), we can set $\omega=d/\lambda_{\min}(\bm{Q})=d^{2}$. Substituting these into our surrogate regret upper bound in (8), we obtain
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2^{5/3}\left(dBC_{x}T\right)^{2/3}.$
@@
Thus, $\lambda_{\min}(\bm{Q})=\min_{\|\bm{x}\|_{2}=1}\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]\geq\frac{m(d-m)}{d^{2}}$ holds, where the equality is from Cesa-Bianchi and Lugosi [8, Lemma 2]. Since we have $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, from (D.4), we can set $\omega=\frac{d^{5}}{4m(d-m)}\geq\|\bm{V}^{-1}\|_{\mathrm{F}}^{2}/\lambda_{\min}(\bm{Q})$. Therefore, our surrogate regret upper bound in (8) is reduced to
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2\left(\frac{d^{5}}{m(d-m)}\right)^{1/3}(BC_{x}T)^{2/3}.$
@@
We finally consider online ranking with the Hamming loss and the number of items $m$. From Cesa-Bianchi and Lugosi [8, Proposition 4], the smallest positive eigenvalue is at least $1/m$. Hence, since $\bm{V}=-\frac{1}{m}\bm{I}$, we have
@@
where we used $\mathrm{rank}(\bm{Q}^{+})\leq d=m^{2}$, and this allows us to choose $\omega=m^{5}$. Substituting these values into our surrogate regret upper bound in (8) , we obtain
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+(2m)^{5/3}(BC_{x}T)^{2/3}.$
Appendix E Details omitted from Section 4
@@ E.1 Details of Optimistic Delayed Adaptive FTRL (ODAFTRL)
We provide a more detailed explanation of the Optimistic Delayed Adaptive FTRL (ODAFTRL) algorithm used for updating $\bm{W}_{t}$ in Section 4. Recall that ODAFTRL computes $\bm{W}_{t}$ by the following update rule:
@@
where $\lambda_{t}\geq 0$ is the regularization parameter. Note that we use the notation $a_{1:t}=\sum_{i=1}^{t}a_{i}$ for simplicity in the following. The ODAFTRL algorithm, when using the parameter update called AdaHedgeD, satisfies the following lemma:
Lemma E.1 ([17, Theorem 12]).
Fix $\alpha>0$. Let $f_{t}\colon\mathcal{W}\to\mathbb{R}$ be a convex function for each $t=1,\dots,T$. Suppose that we update $\lambda_{t}$ in (10) by the following AdaHedgeD update:
@@
$\delta_{t}=\min\big{\{}F_{t+1}(\bm{W}_{t})-F_{t+1}(\bar{\bm{W}}_{t}),\langle\bm{G}_{t},\bm{W}_{t}-\bar{\bm{W}}_{t}\rangle,F_{t+1}(\widehat{\bm{W}}_{t})-F_{t+1}(\bar{\bm{W}}_{t})+\langle\bm{G}_{t},\bm{W}_{t}-\widehat{\bm{W}}_{t}\rangle\big{\}}_{+},$
@@
$\sum_{t=1}^{T}{f_{t}(\bm{W}_{t})}-\sum_{t=1}^{T}{f_{t}(\bm{U})}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq\left(\frac{B^{2}}{2\alpha}+1\right)\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+2\alpha b_{t}}\right),$
@@
In the following, we let $\alpha=\frac{B^{2}}{2}$ for simplicity. Then, since $S_{t}$ is the convex function ,we have
$\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}-\sum_{t=1}^{T}{S_{t}(\bm{U})}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq 2\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+B^{2}b_{t}}\right).$
@@
We present Theorem 4.3 in a more detailed form and provide its proof. In what follows, let
$C_{y}=\max\left\{\max_{\bm{y}\in\operatorname{conv}(\mathcal{Y})}\|\bm{y}\|_{2},\max_{\bm{y},\bm{y}^{\prime}\in\operatorname{conv}(\mathcal{Y})}\|\bm{y}-\bm{y}^{\prime}\|_{2}\right\}$
denote the maximum of the largest Euclidean norm of vectors in $\operatorname{conv}(\mathcal{Y})$ or the diameter of $\operatorname{conv}(\mathcal{Y})$.
@@
Let $\alpha=\frac{B^{2}}{2}$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a fixed delay of $D$ achieves
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BC_{x}C_{y}D+\frac{2bB^{2}}{a}+\min\left\{\frac{b(D+1)^{2}}{2},\frac{3}{2}\left(a^{-1}bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\right)^{1/3}\right\}.$
@@
From the definition of $b_{t}$, it holds that
$b_{t}$ $\leq\min\left\{\frac{1}{2}\|\bm{G}_{t-D:t}\|_{\mathrm{F}}^{2},\|\bm{G}_{t-D:t}\|_{\mathrm{F}}\|\bm{G}_{t}\|_{\mathrm{F}}\right\}$
$\leq\min\left\{\frac{b(D+1)}{2}\sum_{s=t-D}^{t}S_{t}(\bm{W}_{t}),C_{x}C_{y}(D+1)\sqrt{bT\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}\right\},$
where we used $\|\sum_{i=1}^{n}\bm{A}_{i}\|_{\mathrm{F}}^{2}\leq n\sum_{i=1}^{n}\|\bm{A}_{i}\|_{\mathrm{F}}^{2}$ for any matrix $\bm{A}_{i}$, the Cauchy–Schwarz inequality, $\|\bm{G}_{t}\|_{\mathrm{F}}\leq\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|_{2}\|\bm{x}_{t}\|_{2}\leq C_{x}C_{y}$, and $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$. Combining this inequality with Lemma E.1 and the definition of $a_{t}$, we have
$\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}$
$\leq 4B\max_{s\in[T]}\sum_{i=s-D}^{s-1}\|\bm{G}_{i}\|_{\mathrm{F}}$
$\qquad+2B\sqrt{\sum_{t=1}^{T}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+\min\left\{\frac{b(D+1)}{2}\sum_{t=1}^{T}\sum_{s=t-D}^{t}S_{t}(\bm{W}_{t}),C_{x}C_{y}(D+1)\sqrt{bT\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}\right\}}$
$\leq 4BC_{x}C_{y}D+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
$\qquad+2\min\left\{\sqrt{\frac{b(D+1)^{2}}{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}},\left(bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right\},$
where we used $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$ and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$ in the last inequality. From this inequality and Assumption 4.1, we can the evaluate surrogate regret as
@@
$\leq 4BC_{x}C_{y}D+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
$+2\min\left\{\sqrt{\frac{b(D+1)^{2}}{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}},\left(bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right\}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
$\leq 4BC_{x}C_{y}D+\frac{2bB^{2}}{a}+\min\left\{\frac{b(D+1)^{2}}{a},\frac{3}{2}\left(a^{-1}bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\right)^{1/3}\right\},$
where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq({3}/{4})\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$, which hold for any $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$, in the last inequality. ∎
E.3 Proof of high probability bound
We present the result of the high probability bound in a more detailed form and provide its proof.